OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (tcp2315):

let x be the larger number, and let y be the smaller number x+y=170 x-y=26 rearrange teh second for x... x=26+y then sub it in (26+y)+y=170 2y=144 y=77 x-77=26 x=103